Write the equation of the line which passes through (2, 1) and is perpendicular to x = –2.
Answer by stanbon(75887)

Write the equation of the line which passes through (2, 1) and is perpendicular to x = –2.
------------
x = -2 is a vertical line with an undefined slope.
Any line perpendicular to it must be a horizontal
line with slope = 0.
--------------------------
Equation of line thru (2,1) with slope = 0
Form: y = mx+b
1 = 0*2 + b
b = 1
---
Equation:
y = 0*x + 1
y = 1
===============
Cheers,
Stan H.